# Density estimation via mixture discrepancy and moments

### v1

@@ 5 Conclusion and discussion
The main contributions of this work are to propose the DSP-mix and MSP methods. They still adopt the framework of the DSP method in [4] to learn a adaptive piecewise constant function to approximate the underlying probability density but respectively use mixture discrepancy and moments instead of star discrepancy as the uniformity measurements for observations. The mixture discrepancy and moments are easier to compute and theoretically maintain reflection invariance and rotation invariance, which are the shortcomings of the star discrepancy. Numerical experiments indicate that DSP-mix and MSP run approximately ten times faster than DSP while maintaining the same accuracy. There are still some issues worthy of further research, such as the quantitative error analysis of DSP-mix and MSP, the application to practical density estimation problems in higher dimensions, and more efficient uniformity measurement approaches.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 12325112, 12288101) and the High-performance Computing Platform of Peking University.

### v2

@@ 5 Conclusion and discussion
This work proposes the DSP-mix and MSP methods, both of which adopt the same framework of the DSP method in [18] to learn a adaptive piecewise constant function to approximate the underlying probability density. The novelty lies in the measures of uniformity for observations. DSP-mix and MSP adopt the mixture discrepancy and moments, respectively, instead of the star discrepancy in DSP. The mixture discrepancy and moments are easier to compute and theoretically maintain reflection invariance and rotation invariance, which are the shortcomings of the star discrepancy. Numerical experiments up to 30 dimension demonstrate that MSP can maintain the same accuracy compared with DSP, while gaining an increase in speed by a factor of two to twenty for large sample size. DSP-mix achieves satisfactory accuracy and boosts the efficiency in low-dimensional tests ($d\leq 6$), but might lose accuracy in high-dimensional problems due to a reduction in partition level.
As a summary, the accuracy of DSP, DSP-mix and MSP can be improved by increasing the sample size $N$ under appropriate partition level $L$, although it still suffers from the overfitting problem when $L$ is too large. Fortunately, the overfitting can be alleviated as more samples are used. In practice, we can utilize the trend $L\sim\theta^{-1}$ to estimate the partition level by first testing on a small dataset and using extrapolation. This gives us a simple criterion to control the computational complexity.
There are still some issues worthy of further research, such as the quantitative error analysis of DSP-mix and MSP, the application to practical density estimation problems in higher dimensions, and more efficient uniformity measurement approaches.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 12325112, 12288101, 12571413) and the High-performance Computing Platform of Peking University. The authors are thankful to the anonymous referee for his/her valuable suggestions.